In the Solow model, suppose the per worker production function is y: 4 k . Suppose s : 009, n-o 07, and d-012. Calculate the steady-state equilibrium capital-labor ratio. k(Round to two decimal places.)
Answer #1

Per worker production function is y = 4k^0.5

Steady state capital labor ratio or capital per woker is kss

kss/yss = s/d +n

k/ 4k^0.5 = 0.09/(0.07 + 0.12)

k^0.5 = (0.09*4)/(0.07 + 0.12)

k = 3.59

This is the required value